OpenStudy (anonymous):

Write a proportion for the following situation. Your car averages 18 miles per gallon on the highway. If the gas costs $1.85 per gallon, how much does it cost in dollars per mile to drive your car on the highway?

OpenStudy (anonymous):

\[x=18\]\[x∗1.85=33.3\]\[33∗(3/10)\]\[99:10\]

OpenStudy (anonymous):

the answer is 99/10?

OpenStudy (anonymous):

If you are looking for the end ratio, yes 1.85 is converted to 185/100 or 37 / 20 So the end ratio is 37/27=99/10

OpenStudy (anonymous):

needs an edit button, I meant 37/20=99/10

OpenStudy (anonymous):

so then that'll be 9 and 9/10 or 9.9?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thnx, bro!!